OpenStudy (anonymous):

Question 5 (Multiple Choice Worth 1 points) If f(x) = 2x − 2 and g(x) = 3x − 1, find (f − g)(4). 17 −17 5 −5

OpenStudy (ghazi):

(3x-1)-(2x-2)= x+1 ...put x=4